It's been a long time since I had to use this stuff. I want to know how to calculate shipping and handling percentages and tax percentages. Here's an example:
$256.50.....with a 40% discount, then
$102.60.....with a 6% shipping and
then I need to figure out a state tax percentage 7.0%
and I need the answer, the whole total and how it was calculated.
thanks

Answer by AdolphousC(70)

An easy way to solve problems like this is to remember that to get your answer you need to multiply your original total by one of the two following quatities...
if you have an additional charge such as tax you would multiple by ( 1 + tax ) for example a 6% tax would be written as ( 1 + .06 ) or (1.06) by multiplying your original price by this (1.06) you would get the new price with the tax included.


a discount is similar, except instead of adding you will subtract, so a 40% discount would be ( 1 - .4 ) or ( 0.6 ), multiplying your total by .6 will give you your discount price


Now, lets look at your question specficially
$256.50 with a 40% discount, so multiply
(256.50) * (1 - .4) simplify and you have (256.50)*(0.6) = $153.90


$102.60 with a 6% shipping cost, so multiply
(102.60)* (1 + .06)simplify and you have (102.60)*(1.06) = $108.76


To get the out od state tax, add the two items together, and multiply by the tax rate using the same method as shipping ( 1 + .07 )
( 153.90 + 108.76 ) * ( 1.07 )
( 262.66 ) * (1.07) = $281.04 ( this is your total )